OpenStudy (anonymous):

(F o G)(x) F(x)=x-6 G(x)=x-1

OpenStudy (anonymous):

get rid of the circle first that is always the first step \[G\circ F(x)=G(F(x))\]

OpenStudy (anonymous):

second step is to replace the general \(G\) by the specific one you have in this case \[F(G(x))=F(x-1)\]

OpenStudy (anonymous):

third step is to replace the \(x\) in \(F(x)\) by \(x-1\) you get \[F(x-1)=(x-1)-6\]

OpenStudy (anonymous):

last step is to do whatever algebra is needed to clean it up in this case all you need to do is write \((x-1)-6-x-7\)

OpenStudy (anonymous):

Well how did you get the -x-7?